### v1

@@ Abstract
With the rapid advancement of artificial intelligence technology, Large Language Models (LLMs) have demonstrated remarkable potential in the field of Natural Language Processing (NLP), including areas such as content generation, human-computer interaction, machine translation, and code generation, among others. However, their widespread deployment has also raised significant safety concerns. In recent years, LLM-generated content has occasionally exhibited unsafe elements like toxicity and bias, particularly in adversarial scenarios, which has garnered extensive attention from both academia and industry. While numerous efforts have been made to evaluate the safety risks associated with LLMs, there remains a lack of systematic reviews summarizing these research endeavors. This survey aims to provide a comprehensive and systematic overview of recent advancements in LLMs safety evaluation, focusing on several key aspects: (1) "Why evaluate" that explores the background of LLMs safety evaluation, how they differ from general LLMs evaluation, and the significance of such evaluation; (2) "What to evaluate" that examines and categorizes existing safety evaluation tasks based on key capabilities, including dimensions such as toxicity, robustness, ethics, bias and fairness, truthfulness, and so on; (3) "Where to evaluate" that summarizes the evaluation metrics, datasets and benchmarks currently used in safety evaluations; (4) "How to evaluate" that reviews existing evaluation toolkit, and categorizing mainstream evaluation methods based on the roles of the evaluators. Finally, we identify the challenges in LLMs safety evaluation and propose potential research directions to promote further advancement in this field. We emphasize the importance of prioritizing LLMs safety evaluation to ensure the safe deployment of these models in real-world applications.
@@ I Introduction
As artificial intelligence technology evolves at an unprecedented pace, Large Language Models (LLMs) have demonstrated remarkable potential across various fields, becoming a central focal point in both the tech industry and academic circles. These sophisticated models have achieved significant breakthroughs in Natural Language Processing (NLP) and exhibit robust capabilities in diverse application scenarios, including content generation[66], human-computer interaction[67], machine translation[64], and code generation[65], among others [233, 235]. LLMs are designed to process and generate human-like text, leveraging the vast amounts of data on which they are trained. This training enables them to produce fluent and coherent text while demonstrating a degree of comprehension and reasoning abilities[68], significantly advancing the progress of artificial intelligence.
@@
However, like any emerging technology, the widespread application of large language models brings numerous challenges and safety risks. Recent studies have demonstrated that LLMs are prone to generating harmful content, such as toxicity[7], bias[23] and false information [234], particularly in adversarial scenarios where these issues are even more pronounced. In addition, attack methods targeting LLMs safety, such as prompt injection attacks and jailbreak attacks [228, 229], have become increasingly sophisticated, aiming to bypass the safety alignment mechanisms between LLMs and humans, thus inducing LLMs to generate unsafe content[69]. This not only negatively impacts user experience but also raises ethical and legal concerns that could be exploited by malicious actors. Therefore, effectively evaluating the safety of LLMs and ensuring their safe deployment has become a pressing issue that requires urgent attention.
@@
The emergence of LLMs safety evaluation stems from a profound recognition and urgent need to address the potential risks posed by these models and their societal impacts. In the processes of LLMs research, development, and deployment, safety evaluation plays a crucial role. Unlike general large language model evaluations[70, 66, 71], LLMs safety assessment not only requires the model to possess fundamental general capabilities, such as language understanding, generation, and reasoning, but also places particular emphasis on the safety issues associated with the content generated by the model. The assessment covers a range of critical dimensions, including toxicity, robustness, morality, bias and fairness, and credibility. By conducting a systematic safety assessment of LLMs, potential risks can be identified and mitigated in a timely manner, thereby ensuring the safety of these models in practical applications. This rigorous evaluation process is essential for enhancing user trust and preventing societal issues that may arise from model deficiencies, ultimately fostering the healthy development and widespread adoption of LLM technologies.
While numerous researchers have actively explored the safety evaluation of large language models in recent years, a systematic summary of the existing body of research remains conspicuously absent. Existing studies either focus on specific issues or evaluation methods or fail to provide a comprehensive classification, organization, and synthesis of safety evaluations[72, 73]. As a result, they fail to offer a holistic view of current methods, challenges, and future directions in LLMs safety evaluation.
@@
Against this backdrop, this paper aims to provide a thorough and systematic review of recent advancements in LLMs safety evaluation, thereby addressing this research gap. Specifically, as shown in Figure 1, we delve into existing work from the following four dimensions: (1) "Why Evaluate" elucidates the background of current LLMs safety evaluation, distinguishing it from general LLM evaluations. This section underscores the significance of safety evaluations in ensuring that LLMs can be responsibly deployed in real-world applications, highlighting the potential risks associated with unassessed models. (2) "What to Evaluate" summarizes the primary tasks associated with LLMs safety evaluation, presenting a detailed classification that encompasses various facets such as toxicity, robustness, ethics, bias and fairness, truthfulness, and more. (3) "Where to Evaluate" compiles current commonly used evaluation metrics and categorizes the datasets and benchmarks employed in the field. The aim of this section is to provide researchers with a comprehensive reference for selecting appropriate evaluation criteria, thereby facilitating faster progress in safety evaluation research. (4) "How to Evaluate" reviews existing evaluation toolkits and categorizes evaluation methods based on the roles of the evaluators, whether they are automated systems or human assessors, thus providing insights into the methodologies and best practices for conducting safety evaluations. Finally, we discuss the current challenges facing the field of LLMs safety evaluation and outline potential future research directions.
@@
  • Comprehensive Review: To the best of our knowledge, this paper presents the first comprehensive and systematic survey of recent advancements in the field of large language model safety evaluation, addressing a significant gap in the existing literature.
  • Clear Classification Framework: We establish a detailed classification framework that delineates the primary tasks of LLMs safety evaluation across various dimensions, enhancing understanding within the research community.
  • Consolidated Evaluation Resources: We compile and categorize current commonly used evaluation metrics, datasets, benchmarks, evaluation toolkits, and methods, offering a comprehensive resource for researchers to reference and apply in safety evaluations.
  • Future Research Directions: We discuss the key challenges currently facing the field and outline potential avenues for future research, thereby providing guidance to promote the healthy development of LLMs safety evaluation.
{forest}
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=root-line-color, rounded corners, align=center, text centered, minimum width=5em, edge+=edge-color, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=1pt, , where level=1 draw=child-one-line-color, text width=13em, font=, , where level=2 draw=child-two-line-color, text width=20em, font=, , where level=3 draw=child-three-line-color, minimum width=28em, font=, , [ LLMs Safety Evaluation, for tree=fill=root-color, [ Why to evaluate §$\S$ II, for tree=fill=child-one-color, [ Background (A), for tree=fill=child-two-color ] [ LLMs evaluation vs
@@
(B), for tree=fill=child-two-color, ] [ Importance (C), for tree=fill=child-two-color, ] ] [ What to evaluate §$\S$ III, for tree=fill=child-one-color, [ Toxicity (A), for tree=fill=child-two-color, [ ToxiGen [2]; ToxiChat [1]
@@
AdvBench [7]; Red Teaming ChatGPT [6], for tree=fill=child-three-color ] ] [ Robustness (B), for tree=fill=child-two-color, [ Chen et al. [53]; DecodingTrust [136]
@@
An et al. [58]; Varshney et al. [48], for tree=fill=child-three-color ] ] [ Ethics (C), for tree=fill=child-two-color, [ TECHHAZARDQA [14]; LOCALVALUEBENCH [15]
TRUSTGPT [16]; MoralBench [20]
CVALUES [46]; Sun et al. [17]
Zhang et al. [105]; Hendrycks et al. [106], for tree=fill=child-three-color ] ] [ Bias and Fairness (D), for tree=fill=child-two-color, [ Cantini et al. [13]; Zhang et al. [23]
@@
Azeem et al. [22]; Zhou et al. [6]
Wang et al. [116], for tree=fill=child-three-color ] ] [ Truthfulness (E), for tree=fill=child-two-color, [ AI-LIEDAR [26]; THaMES [127]
FACTKB [129]; ReEval [146]
HaluEval [147]; Wang et al. [152]
GraphEval [153]; Zhu et al. [157], for tree=fill=child-three-color ] ] [ Other (F), for tree=fill=child-two-color, [ LLM-PBE [39]; Park et al. [34]
MedSafetyBench [183]; CEB [114]
CYBERSECEVAL [36]; CodeSecEval [35]
SHIELD [37]; R-Judge [40]
RiskAwareBench [41]; ASB [42], for tree=fill=child-three-color ] ] ] [ Where to evaluate §$\S$ IV, for tree=fill=child-one-color, [ Metric (A), for tree=fill=child-two-color, [ Cai et al. [177]; StrongReject [184]
@@
Mei et al. [190], for tree=fill=child-three-color ] ] [ Toolkit (B), for tree=fill=child-two-color, [ Sun et al. [17]; HarmBench [10]
JailbreakBench [11]; EasyJailbreak [192]
WalledEval [54]; JailbreakEval [187]
, for tree=fill=child-three-color ] ] [ Datasets and Benchmarks (C), for tree=fill=child-two-color, [ Chu et al. [63]; R-Judge [40]
@@
XSafety [185]; JailJudge OOD [194]
PolyglToxicityPrompts [5], for tree=fill=child-three-color ] ] ] [ How to evaluate §$\S$ V, for tree=fill=child-one-color, [ Human-based Evaluation (A), for tree=fill=child-two-color, [ Meadows et al. [15]; Alghamdi et al. [28]
Movva et al. [209], for tree=fill=child-three-color ] ] [ Rule-based Evaluation (B), for tree=fill=child-two-color, [ Liu et al. [37]; Ye et al. [210], for tree=fill=child-three-color ] ] [ Model-based Evaluation (D), for tree=fill=child-two-color, [ Cai et al. [177]; Sun et al. [17]
@@
JailJudge Guard [194], for tree=fill=child-three-color ] ] ] [ Challenges and
future directions
§$\S$ VI, for tree=fill=child-one-color, [ Unified Evaluation (A), for tree=fill=child-two-color ] [ Dynamic Evaluation (B), for tree=fill=child-two-color ] [ Reliable and Efficient Evaluator (C), for tree=fill=child-two-color ] [ Specific Application Evaluation (D), for tree=fill=child-two-color ] [ Beyond Evaluation: Safe to Responsible (E), for tree=fill=child-two-color ] ] ]
@@ II WHY TO EVALUATE
II-A Background
The extensive adoption of large language models (LLMs) has showcased remarkable utility across diverse domains, including knowledge inference [75, 76], content generation [74], code development support [77, 78], and data extraction [79, 232, 231]. However, the substantial generative capabilities of these models are not without risks. Studies indicate that, when guided by certain prompts, LLMs can produce offensive or prejudiced outputs. Moreover, when faced with topics outside their training scope, these systems often construct seemingly credible yet incorrect information, contributing to misinformation. Such fabricated and misleading content may be exploited with malicious intent, causing significant societal harm [5, 24, 80, 21, 81, 82].
For instance, in the financial sector, LLMs may generate inaccurate credit scores due to biases or deficiencies in training data, resulting in capable borrowers being mistakenly classified as high-risk and denied loans [83]. In the field of smart education, LLMs may perpetuate stereotypes about the learning abilities of certain demographic groups, leading to systematic recommendations of overly simplified learning content and limiting the long-term development of these students [84]. In the healthcare domain, incorrect diagnostic suggestions generated by LLMs could severely impact patient treatment outcomes and even endanger lives [85].
The security challenges of LLMs primarily manifest in four key aspects:
  1. 1.
    Generation of offensive content: LLMs are prone to producing harmful or objectionable language, including discriminatory, racist, or hateful speech [3].These risks are particularly concerning when LLMs are applied in sensitive environments, such as automated customer service or social media platforms.
  2. 2.
    Generation of biased and unethical content: LLMs may perpetuate societal biases embedded within their training data, such as those related to gender, ethnicity, or religion, leading to discriminatory outcomes [86]. If left unaddressed, these biases not only reinforce existing inequalities but may also erode trust in the technology.
  3. 3.
    Generation of factually incorrect content: LLMs may propagate factual inaccuracies or spread rumors, particularly when trained on datasets containing misinformation or widely circulated false news [87]. These models lack the ability to verify the truthfulness of information, and as a result, present inaccuracies with an air of confidence, making it difficult for users to distinguish between valid knowledge and falsehoods.
  4. 4.
    Vulnerability to adversarial attacks: LLMs are susceptible to adversarial manipulations, such as prompt injection and jailbreak attacks, where attackers craft inputs designed to circumvent the model’s safety protocols [9]. These attacks can prompt the model to produce harmful instructions or generate inappropriate content, posing significant security threats.
@@ II-A Background
This dual-edged nature of LLMs underscores the need for comprehensive safeguards and systematic security assessments. Identifying and addressing these risks proactively is crucial to prevent unintended consequences, protect users, and ensure these models contribute positively to society.
@@ II-B LLMs evaluation vs LLMs Safety Evaluation
In this context, LLM safety evaluation has become an indispensable component of the development and deployment process. The objective of safety evaluation extends beyond merely identifying and mitigating potential risks within the model—it aims to ensure that these risks do not adversely impact users or society in real-world applications [88]. Although LLMs safety evaluation shares certain overlaps with conventional model evaluation, their goals and areas of emphasis differ significantly:
  • LLMs Evaluation: This evaluation primarily focuses on assessing the fundamental performance of the model, including its accuracy, fluency, and reasoning capabilities. Standardized datasets and benchmarks are employed to ensure consistency and comparability across tasks. Metrics such as BLEU [89] and ROUGE [90] are commonly used to evaluate the effectiveness of the model in specific applications, such as machine translation [91] and question answering systems [78]. The primary goal of LLM evaluation is to quantify how well the model performs under controlled conditions, emphasizing technical proficiency in language generation.
  • LLMs Safety Evaluation: Building upon the foundation of LLM evaluation, safety evaluation goes further by focusing not only on technical performance but also on the ethical, social, and contextual dimensions of the model’s behavior. It emphasizes ethical integrity, robustness,and bias mitigation to ensure that the model’s outputs align with societal values and remain appropriate across diverse real-world scenarios. In this framework, the ability to generate accurate and fluent text is a necessary but insufficient criterion; the evaluation also examines whether the model produces fair and unbiased outputs, maintains stability under adversarial conditions, and avoids generating toxic or harmful content. In essence, safety evaluation extends beyond the quantitative metrics of traditional assessments, integrating considerations such as the responsible use of AI and the model’s resilience in unpredictable environments.
In contrast to standard performance assessments, LLMs safety evaluation adopts a social responsibility perspective, ensuring that the model’s outputs align with compliance, fairness, and reliability standards across a variety of real-world contexts. While traditional evaluations measure technical proficiency, safety assessments extend this focus by addressing the broader societal impact of the model’s behavior. These evaluations ensure not only that the model performs well but also that it generates content that is ethical, appropriate, and trustworthy in sensitive applications such as healthcare, education, and finance. Key components of safety evaluation include:
  • Toxicity Detection: This process evaluates whether the model produces offensive, harmful, or abusive language during content generation [92]. It aims to prevent the propagation of toxic expressions that could damage user experiences or incite social harm, particularly in public-facing or automated communication systems.
  • Ethics and Bias Evaluation: This assessment identifies and mitigates biases or discriminatory tendencies present in the model’s outputs, such as those targeting specific demographic groups based on gender, ethnicity, or socio-economic status [93, 22].
  • Truthfulness and Credibility: This component ensures that the model generates factually accurate information, minimizing the risk of misleading users [30, 29]. In high-stakes domains like healthcare or legal advice, credibility is essential for maintaining trust and enabling reliable decision-making
  • Robustness Testing: This examines the model’s ability to maintain safe and stable behavior when exposed to adversarial inputs or unexpected scenarios [13, 12]. Robustness evaluation ensures that the model can withstand attempts to manipulate its outputs or bypass safety mechanisms, helping to safeguard against misuse.
Similar to LLMs evaluation, the results of LLMs safety Evaluation provide critical insights that guide model improvement across its design, training, and deployment phases. However, safety evaluation plays a more proactive role by identifying vulnerabilities and addressing them before deployment to prevent unintended harm. Developers can leverage various strategies to enhance model safety, such as reinforcement learning with human feedback (RLHF) to improve the model’s ability to filter toxic content [94], adversarial training to strengthen the model’s defenses against manipulative attacks [95], and enhancing the quality of pre-training data to minimize hallucinations and improve factual accuracy [96]. Ultimately, these efforts are essential for building robust, responsible, and trustworthy AI systems, reducing risks, and ensuring that the outputs align with ethical and societal standards.

### v2

@@ Abstract
With the rapid advancement of artificial intelligence, Large Language Models (LLMs) have shown remarkable capabilities in Natural Language Processing (NLP), including content generation, human-computer interaction, machine translation, and code generation. However, their widespread deployment has also raised significant safety concerns. In particular, LLM-generated content can exhibit unsafe behaviors such as toxicity, bias, or misinformation, especially in adversarial contexts, which has attracted increasing attention from both academia and industry. Although numerous studies have attempted to evaluate these risks, a comprehensive and systematic survey on safety evaluation of LLMs is still lacking. This work aims to fill this gap by presenting a structured overview of recent advances in safety evaluation of LLMs. Specifically, we propose a four-dimensional taxonomy: (i) Why to evaluate, which explores the background of safety evaluation of LLMs, how they differ from general LLMs evaluation, and the significance of such evaluation; (ii) What to evaluate, which examines and categorizes existing safety evaluation tasks based on key capabilities, including dimensions such as toxicity, robustness, ethics, bias and fairness, truthfulness, and related aspects; (iii) Where to evaluate, which summarizes the evaluation metrics, datasets and benchmarks currently used in safety evaluations; (iv) How to evaluate, which reviews existing mainstream evaluation methods based on the roles of the evaluators and some evaluation frameworks that integrate the entire evaluation pipeline. Finally, we identify the challenges in safety evaluation of LLMs and propose promising research directions to promote further advancement in this field. We emphasize the necessity of prioritizing safety evaluation to ensure the reliable and responsible deployment of LLMs in real-world applications.
@@ I Introduction
As artificial intelligence continues to advance at an unprecedented pace, Large Language Models (LLMs) have rapidly emerged as a focal point in both academia and industry. These models have achieved significant breakthroughs in Natural Language Processing (NLP) and demonstrated strong capabilities across diverse applications, including content generation [1], human-computer interaction [2], machine translation [3], and code generation [4], among others [5, 6]. By leveraging massive training corpora, LLMs can generate coherent, human-like text while exhibiting a degree of comprehension and reasoning ability [7], thus driving rapid progress in the broader field of artificial intelligence.
Figure 1: Overview of the Comprehensive and Systematic Survey on Safety Evaluation of LLMs.
@@
However, like any emerging technology, the widespread application of LLMs has introduced critical safety challenges. Recent studies have demonstrated that LLMs are prone to generating harmful content, such as toxicity[8], bias[9] and false information [10], with such issues often exacerbated under adversarial scenarios. In addition, attack methods targeting LLMs safety, such as prompt injection attacks and jailbreak attacks [11, 12], have become increasingly sophisticated, seeking to bypass alignment mechanisms and induce unsafe behavior from LLMs [13]. These vulnerabilities not only degrade user experience but also raise profound ethical and legal risks, offering opportunities for malicious exploitation. Consequently, ensuring the safety of LLMs through systematic evaluation has become an urgent priority.
Table I: A Comparative Analysis of Related Surveys and Studies on Safety Evaluations of LLMs
Papers Year Key Objectives of Safety Evaluation
Taxonomy of Tasks Metrics Datasets and Benchmarks Taxonomy of Evaluators Frameworks Challenges and Future Directions
Guo et al. [14] 2023
Dong et al. [15] 2024
Yi et al. [16] 2024
Ran et al. [17] 2024
Röttger et al. [18] 2025
Ours 2025
@@
Safety evaluation of LLMs is the systematic process of examining whether a model can resist harmful instructions, avoid generating unsafe outputs, and maintain robustness under both benign and adversarial interactions, thereby assessing the degree of risk that LLMs may pose to social order, public interest, and human well-being. The emergence of this research area stems from a growing recognition and urgent need to address the potential risks posed by these models and their societal impacts. Within the broader lifecycle of LLMs research, development, and deployment, safety evaluation plays a crucial role. Unlike general ability evaluation of LLMs [14, 1, 19], which focus primarily on general capabilities such as language understanding, generation, and reasoning, safety evaluation places particular emphasis on the risks and ethical concerns associated with model outputs. The evaluation covers a range of critical dimensions, including toxicity, robustness, morality, bias and fairness, and credibility. By conducting a systematic safety evaluation of LLMs, potential risks can be identified and mitigated in a timely manner, thereby ensuring the safety of these models in practical applications. This rigorous evaluation process is essential for enhancing user trust and preventing societal issues that may arise from model deficiencies, ultimately fostering the healthy development and widespread adoption of LLMs technologies.
While numerous researchers have actively explored the safety evaluation of LLMs in recent years, the field still lacks a comprehensive and systematic review. As shown in Table I, most existing relevant surveys concentrate on general ability evaluation of LLMs [14] on specific topics such as jailbreak attacks [15, 16], with only limited coverage of safety evaluation, failing to form a systematic classification framework and content synthesis. Furthermore, existing studies often remain fragmented, addressing single aspects such as datasets and benchmarks [18] or evaluators [17], lacking a holistic analysis of the entire pipeline, which includes evaluation methods, technical challenges, and future directions. This fragmentation underscores the urgent need for a structured and integrative survey on safety evaluation of LLMs.
@@
Against this backdrop, this survey aims to provide a comprehensive and systematic review of recent advancements in safety evaluation of LLMs, thereby addressing this research gap. Specifically, as shown in Figure 1, we delve into existing work from the following four dimensions: (i) Why to Evaluate, which elucidates the background of safety evaluation, distinguishing it from general ability evaluation. This section underscores the significance of safety evaluations in ensuring that LLMs can be responsibly deployed in real-world applications, highlighting the potential risks associated with unassessed models; (ii) What to Evaluate, which summarizes the primary tasks associated with safety evaluation of LLMs, presenting a detailed classification that encompasses various facets such as toxicity, robustness, ethics, bias and fairness, truthfulness, and more. (iii) Where to Evaluate, which compiles current commonly used evaluation metrics, and categorizes the datasets and benchmarks employed in the field. The aim of this section is to provide researchers with a comprehensive reference for selecting appropriate evaluation criteria, thereby facilitating faster progress in safety evaluation research; and (iv) How to Evaluate, categorizes evaluation methods based on the roles of the evaluators, whether they are automated or manual. It also reviews several evaluation frameworks, providing insights into the methodologies and best practices for conducting safety evaluations. Finally, we highlight the key challenges in this field and outline promising avenues for future research.
@@
  • Comprehensive Review: To the best of our knowledge, this is the first survey to present a systematic and holistic synthesis of recent advancements in safety evaluation of LLMs, addressing a significant gap in the literature.
  • Clear Classification Framework: We establish a detailed taxonomy that delineates the primary tasks of safety evaluation across multiple dimensions, providing conceptual clarity for the research community.
  • Consolidated Evaluation Resources: We compile and categorize existing evaluation metrics, datasets, benchmarks, and methodologies, offering a consolidated reference for future research and practice.
  • Future Research Directions: We highlight key challenges currently facing the field and propose potential research directions to promote the responsible and sustainable development of safety evaluation.
[…]
@@
Why to evaluate $\S$ II Background and Foundations (A) LLMs Evaluation vs LLMs Safety Evaluation (B) Importance (C) What to evaluate $\S$ III Toxicity (A)
ToxiGen [20]; ToxiChat [21]
@@
zhuo et al. [26]
Robustness (B)
Chen et al. [27]; DecodingTrust [28]
@@
An et al. [33]; Varshney et al. [34]
Ethics (C)
TechHazardQA [35]; LocalValueBench [36]
MoralBench [37]; CVALUES [38]
Sun et al. [39]; Hendrycks et al. [40]
Bias and Fairness (D)
Cantini et al. [41]; Zhang et al. [9]
@@
Azeem et al. [45]; Wang et al. [46]
Leidinger et al. [47]
Truthfulness (E)
AI-LIEDAR [48]; THaMES [49]
ReEval [50]; HaluEval [51]
Iqbal et al. [52]; GraphEval [53]
KG-FPQ et al. [54]
Other (F)
LLM-PBE [55]; MedSafetyBench [56]
CYBERSECEVAL [57]; CWEVAL [58]
SafeGenBench [59]; SHIELD [60]
Trident-Bench [61]
Where to evaluate $\S$ IV Metrics (A)
Cai et al. [62]; StrongReject [63]
@@
Mei et al. [64]
Datasets and Benchmarks (B)
Chu et al. [65]; R-Judge [66]
@@
AnswerCarefully [85]; XSafety [86]
JailJudge OOD [69]; PolyglToxicityPrompts [23]
How to evaluate $\S$ V Human-based Evaluation (A)
Meadows et al. [36]; Movva et al. [87]
Souly et al. [63]
Rule-based Evaluation (B)
Zou et al. [8]; Liu et al. [88]
Zeng et al. [89]; Liu et al. [60]
Model-based Evaluation (C)
OpenAI Moderation API [90]; Perspective API [91]
Azure AI Content Safety API [92]; Wang et al. [67]
Yu et al. [93]; GPT-4 Judge [94]
PAIR [95]; PICT et al. [96]
@@
JailJudge Guard [69]
Evaluation Frameworks (D)
HarmBench [31]; JailbreakBench [30]
EasyJailbreak [103]; WalledEval [104]
JailbreakEval [17]; AISafetyLab [105]
PandaGuard [106]
Challenges and future directions $\S$ VI Unified Evaluation (A) Dynamic Evaluation (B) Reliable and Efficient Evaluator (C) Application-Specific Evaluation (D) Safety Evaluation for MLLMs and Agents (E) Beyond Evaluation: From Safe to Responsible (F)
@@ II WHY TO EVALUATE
This section introduces the background and key concepts of safety evaluation of LLMs, clarifies the distinctions between general evaluation and safety-specific evaluation, and highlights its importance in real-world applications. The goal is to underscore the necessity of safety evaluation and establish a foundation for subsequent discussions on advanced methodologies and practices.
II-A Background and Foundations
LLMs have been rapidly adopted across a wide range of tasks — from knowledge inference and information extraction to content and code generation — due to their strong generative and reasoning capabilities. [107, 108, 109, 110, 111, 112, 113, 114] While these capabilities enable new applications, they also expose concrete and repeatable failure modes that raise safety concerns in practice.
Empirical studies have pointed out several failure modes. LLMs can generate toxic or abusive language under certain prompts, undermining user safety in public-facing systems [115, 22]. They can reproduce or amplify social biases and unethical stereotypes present in training data, yielding discriminatory outputs in downstream decisions [116, 117]. LLMs also exhibit hallucinations — confidently stated but factually incorrect assertions — which degrade trust and can mislead users [118, 119]. Finally, LLMs are susceptible to adversarial inputs, such as prompt injection or jailbreaks, which exploit vulnerabilities in the model’s behavior [81, 93].
These technical failure modes have concrete, domain-specific consequences. In finance, biased or incorrect model outputs can distort credit assessments or automated advice [120]. In education, stereotype-aligned recommendations may narrow learning opportunities [121]. In healthcare, hallucinated or misleading suggestions risk patient safety and clinical misjudgments [122]. Moreover, adversaries can weaponize generation capabilities to scale misinformation or targeted manipulation, amplifying societal harms [16].
@@ II-A Background and Foundations
This dual-edged nature of LLMs underscores the need for comprehensive safeguards and systematic security assessments. Identifying and addressing these risks proactively is crucial to prevent unintended consequences, protect users, and ensure these models contribute positively to society. To achieve this, it is necessary to define a formal framework for safety evaluation, which we describe as the process of assessing whether a model can resist harmful instructions, avoid unsafe outputs, and maintain robustness under both benign and adversarial interactions. Formally:
$R=E(T,D,P,M)$ (1)
where $R$ denotes the outcome of the evaluation , which may be quantitative (e.g., scores) or qualitative (e.g., categorical assessment); $E(\cdot)$ represents the safety evaluation function, mapping the selected tasks, data, evaluation process, and metrics to a final result; $T$ indicates the evaluation tasks that define the targeted safety dimensions. This step answers the question of what to evaluate; $D$ represents the datasets or benchmarks used for testing. This step determines where to evaluate; $P$ corresponds to the evaluation process. This aligns with how to evaluate; and $M$ refers to the metrics that quantify safety performance.
@@ II-B LLMs Evaluation vs LLMs Safety Evaluation
The distinction between LLMs evaluation and LLMs safety evaluation lies not in superficial differences of scope or methodology, but in their fundamentally divergent goals and perspectives. Conventional evaluation addresses the question of “what the model can do”, while safety evaluation is concerned with “what the model should not do”. This asymmetry gives rise to significant differences in their objectives, approaches, and challenges.
First, the two forms of evaluation pursue distinct core objectives. LLMs evaluation is driven by an optimization mindset: its purpose is to measure the upper bound of a model’s capabilities by assessing performance on general tasks such as knowledge, reasoning, or text generation. In contrast, safety evaluation adopts a control mindset: its purpose is to probe the lower bound of risk by examining the model’s behavior under malicious or unexpected conditions. The ultimate goal is not to maximize capability, but to minimize harm, ensuring that the model remains robust, compliant, and trustworthy across diverse real-world applications [123]. This divergence in purpose fundamentally shapes the methodology of each evaluation paradigm.
Second, safety evaluation can be understood as an extension rather than a replacement of capability evaluation. While technical proficiency remains a prerequisite, it is insufficient on its own. A model that excels in accuracy or fluency but fails to avoid harmful, biased, or manipulative outputs cannot be considered safe for deployment. Thus, safety evaluation builds on the foundation of general evaluation by expanding the focus from technical correctness to include ethical and societal considerations. In other words, it asks not only whether the model can perform a task, but also whether it should perform it in a given way.
Third, the quantification of evaluation outcomes reflects a striking asymmetry between the two. Capability evaluation typically relies on objective, mathematically defined indicators that enable standardized comparison across tasks and systems. Metrics like accuracy, F1-score, BLEU [124], and ROUGE [125] are designed to quantify improvements in effective and reliability. By contrast, safety evaluation must grapple with inherently subjective and context-dependent criteria, such as fairness, harm, or appropriateness. Defining what constitutes "toxic" or "biased" content is a challenging endeavor, as such determinations may vary across cultural and social contexts—factors that render universal standardization elusive. Presently, there remains a lack of consensus within the research community regarding how to formulate precise definitions of such content [17]. This subjectivity poses unique challenges for designing safety benchmarks and metrics that are both rigorous and widely applicable.
In summary, while both LLMs evaluation and safety evaluation are indispensable, they differ in orientation: one is performance-driven, seeking optimization, while the other is risk-driven, seeking control. The interplay between these two dimensions—capability and safety—defines the trajectory of responsible LLM development, ensuring that models are not only powerful but also aligned with societal values and constraints.